# A Novel Terabit Grid-of-Beam Optical Wireless Multi-User Access Network With Beam Clustering

### v1

@@ Abstract
In this paper, we put forward a proof of concept for sixth generation (6G) Terabit infrared (IR) laser-based indoor optical wireless networks. We propose a novel double-tier access point (AP) architecture based on an array of arrays of vertical cavity surface emitting lasers (VCSELs) to provide a seamless grid-of-beam coverage with multi-Gb/s per beam. We present systematic design and thorough analytical modeling of the AP architecture, which are then applied to downlink system modeling using non-imaging angle diversity receivers (ADRs). We propose static beam clustering with coordinated multi-beam joint transmission (CoMB-JT) for network interference management and devise various clustering strategies to address inter-beam interference (IBI) and inter-cluster interference (ICI). Non-orthogonal multiple access (NOMA) and orthogonal frequency division multiple access (OFDMA) schemes are also adopted to handle intra-cluster interference, and the resulting signal-to-interference-plus-noise ratio (SINR) and achievable data rate are derived. The network performance is studied in terms of spatial distributions and statistics of the downlink SINR and data rate through extensive computer simulations. The results demonstrate that data rates up to $15$ Gb/s are achieved within the coverage area and a properly devised clustering strikes a balance between the sum rate and fairness depending on the number of users.
@@ I Introduction
Optical wireless communication (OWC) technology based on narrow-beam infrared (IR) lasers has demonstrated the potential to provide ultra-high-speed transmissions for indoor multi-user wireless networks [3, 4, 5, 6, 7]. Beam-steered IR light communication is tailored to provide multi-Gb/s data rates per user by directing narrow IR laser beams at mobile devices [3, 8, 9]. One of the main challenges for the realization of indoor narrow-beam IR systems is beam steering as addressed in published research [10, 5, 6]. In [10], Koonen et al. have proposed and investigated a technique for two-dimensional (2D) steering of directional IR pencil beams (i.e., beams with very low divergence) by means of a passive diffractive optical module composed of crossed grating elements whose output beam angle is controlled by the wavelength of the laser beam. The authors have demonstrated downstream delivery with data rates of $32$ Gb/s and $42.8$ Gb/s per beam using $4$- pulse amplitude modulation (PAM) and adaptive discrete multitone (DMT) modulation, respectively, over $3$ m reach for a wavelength of $1550$ nm. In [5], Koonen et al. have proposed an alternative approach to implement wavelength-controlled 2D steering of IR pencil beams by using a high port-count arrayed waveguide grating router (AWGR) encompassing a bundle of single mode fibers with distinct wavelengths. To this end, they have rearranged the output fibers of the AWGR as a 2D square array and placed a lens next to the array, by which the beam direction for each fiber is determined by varying the position of the array with respect to the lens. The authors have shown $112$ Gb/s $4$-PAM transmission with $2.5$ m reach based on an $80$-ports C-band AWGR, estimating a total throughput of more than $8.9$ Tb/s when all the beams are in use. In [6], Wang et al. have designed a silicon photonic integrated phased array as an optical beam steering device, whereby they have achieved $12.5$ Gb/s error-free transmission over a distance of $1.4$ m using on-off keying (OOK) modulation.
@@
The aforementioned works apply advanced optical designs for beam steering to provide IR laser-class indoor coverage. However, the high complexity and expensive components associated with beam steering systems do not suit large-scale indoor network deployments. Taking a different approach, in [11, 12], Sun et al. have proposed a beam-domain massive multiple input multiple output (MIMO) optical wireless communication (OWC) system in which a large transmit lens is employed in front of a light-emitting diode (LED) array. The lens refracts lights emitted from the LED array elements toward different directions, enabling the optical access point (AP) to simultaneously communicate with a large number of user terminals in a wide coverage area. The authors have presented performance analysis for the optical massive multi-user MIMO transmission system with linear precoding and derived the optimal transmit covariance matrix under the total and per LED power constraints. Based on their findings, as the number of LEDs approaches infinity, the asymptotically optimal multi-user transmission policy under both power constraints is to use non-overlapping light beams.
[…]
@@
In this paper, we propose a novel double-tier AP architecture incorporating an array of VCSEL arrays, which we also refer to as array of arrays for brevity. The AP architecture comprises multiple transmitter elements with each one having a predefined orientation to cover a specific area of the network. Each transmitter element is composed of a VCSEL array and a plano-convex lens. This design offers a simple and scalable solution suitable for establishing an indoor network of densely deployed laser beams to provide a full grid-of-beam coverage through a single AP without the need for beam steering. To unlock the ultra-high capacity of the proposed AP and to eliminate strict alignment requirements, we employ angle diversity receivers using non-imaging concentrators coupled with high-speed photodiode (PD) arrays [17]. Besides, the contributions of this paper are summarized as follows:
@@
  • The multi-user network performance is analyzed in terms of the received signal-to-interference-plus-noise ratio (SINR) and achievable rate for both non-orthogonal multiple access (NOMA) and orthogonal frequency division multiple access (OFDMA) schemes.
  • Various insights are provided into the system performance through extensive computer simulation results, which can be used as design guidelines for the development of laser-based optical wireless access networks.
@@ II Gaussian Beam Propagation
[…]
@@
where $P_{\mathrm{t}}$ is the transmit optical power, and $r$ and $z$ are the radial and axial positions, respectively. The radius of the beam spot at distance $z$ from the beam waist is given by [18]:
@@
By using (2), the divergence angle of the beam in the far field is obtained as [18]:
@@
A Gaussian beam can also be fully characterized based on its complex $q$-parameter, which is defined as [18]:
@@
In fact, the real and imaginary parts of the reciprocal of $q(z)$ encompass $R(z)$ and $w(z)$ as follows:
@@ III-A Array of Arrays Architecture
Fig. 1 depicts an indoor optical wireless network with full beam coverage by deploying a massive number of laser beams based on an array of arrays of VCSELs. The proposed double-tier AP design is composed of $9$ identical transmitter elements arranged as a $3\times 3$ array with a $5\times 5$ VCSEL array incorporated in each element. The total number of VCSELs used in this array of arrays architecture is $N_{\mathrm{VCSEL}}=225$. By choosing appropriate orientation angles for transmitter elements, each VCSEL array covers almost a ninth of the network area, as shown in Fig. 1. Note that the middle element of the AP that covers the central region of the network is not titled.
Figs. LABEL:sub@Fig:AP_a and LABEL:sub@Fig:AP_b show the optical design of the double-tier AP. The adjacent transmitter elements are placed at a center-to-center distance of $d_{\mathrm{lens}}$. In each array, the VCSELs are placed close to each other with a pitch distance of $\delta$. As a result, the beam spots may significantly overlap on the receiver plane. The beams can be separated from one another by using a plano-convex lens in front of each VCSEL array. This allows the light beams to be to refracted towards different directions so as to provide a high spatial resolution on the receiver plane. The refraction angle experienced by each beam depends on the lens thickness and the path it travels through within the lens. In order to achieve the desired spatial resolution, the lens parameters need to be properly adjusted based on the design requirements. The adjustable parameters of the plano-convex lens include the diameter $L$, the radius of curvature $R_{\mathrm{lens}}$, and the center thickness $\tau_{\mathrm{c}}$.
@@ III-B1 Light Refraction
When light traverses a path between two media, it is refracted at the boundary. According to Snell’s law [18]:
@@
where $n_{1}$ and $\theta_{1}$, respectively, are the refractive index and the incident angle in the first medium, and $n_{2}$ and $\theta_{2}$ are the corresponding parameters for the second medium. Snell’s law can be expressed in vector form, leading to the law of refraction [18]:
$\mathbf{n}\times\mathbf{v}_{1}=\mu(\mathbf{n}\times\mathbf{v}_{2}),\quad\mu=\frac{n_{1}}{n_{2}},$ (8)
where $\mathbf{n}$ is the unit vector normal to the boundary surface, $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ are the normalized incidence and refraction vectors, respectively, and $\times$ denotes the cross product. The refraction vector is given by [18]:
@@ III-B2 Beam Transformation
In order to derive the parameters of the transformed beam, we use the $q$-parameter transformation given by [18, 20]:
@@
By substituting the ABCD parameters from (III-B2) into (21), the distance of the transformed beam waist from the tangent plane on the lens surface and the Rayleigh range of the transformed beam are derived as:
@@
where the operators $\mathrm{Re}$ and $\mathrm{Im}$ take the real and imaginary parts of a complex variable, respectively, and $f$ is the effective focal distance of the lens defined as [18]:
@@ III-C Total Spatial Intensity Distribution
In order to derive the total spatial intensity of the AP, the transmitter elements are indexed according to Fig. LABEL:sub@Fig:AP_b. Based on Euler angles with clockwise rotations, the titling of the $v$th transmitter element about the $x$ and $y$ axes can be modeled using the following rotation matrices:
$\mathbf{R}_{x}(\alpha_{v})$ $=\left[\begin{matrix}1&0&0\\ 0&\cos\alpha_{v}&-\sin\alpha_{v}\\ 0&\sin\alpha_{v}&\cos\alpha_{v}\end{matrix}\right],$ (26d)
$\mathbf{R}_{y}(\beta_{v})$ $=\left[\begin{matrix}\cos\beta_{v}&0&\sin\beta_{v}\\ 0&1&0\\ -\sin\beta_{v}&0&\cos\beta_{v}\end{matrix}\right],$ (26h)
@@
where $\mathbf{v}^{\prime}_{i}$ is evaluated for $i=j-25(v-1)$ based on (16). The coordinates of the $v$th transmitter element are given by:
$\mathbf{q}_{v}=\left[\begin{matrix}x_{v}\\ y_{v}\\ z_{v}\end{matrix}\right]=\mathbf{R}_{y}(\pi)\left[\begin{matrix}x^{\prime}_{v}\\ y^{\prime}_{v}\\ d_{\mathrm{c}}\end{matrix}\right]+\left[\begin{matrix}0\\ 0\\ h_{\mathrm{DL}}+d_{\mathrm{c}}\end{matrix}\right]=\left[\begin{matrix}-x^{\prime}_{v}\\ y^{\prime}_{v}\\ h_{\mathrm{DL}}\end{matrix}\right],$ (29)
where $d_{\mathrm{c}}=d_{\mathrm{VL}}+\tau_{\mathrm{c}}$, and $\left[x^{\prime}_{v}~{}y^{\prime}_{v}\right]^{T}$ is equal to the $v$th column of $\mathbf{Q}_{\mathrm{t}}\in\mathbb{R}^{2\times 9}$ given by:
@@
Refer to caption (a) $R_{\mathrm{lens}}=25$ mm Refer to caption (b) $R_{\mathrm{lens}}=17$ mm Refer to caption (c) $R_{\mathrm{lens}}=15$ mm
Refer to caption (a) $R_{\mathrm{lens}}=15$ mm Refer to caption (b) $R_{\mathrm{lens}}=13$ mm Refer to caption (c) $R_{\mathrm{lens}}=10$ mm

### v2

@@ Abstract
In this paper, we put forward a proof of concept for sixth generation (6G) Terabit infrared (IR) laser-based indoor optical wireless networks. We propose a novel double-tier access point (AP) architecture based on an array of arrays of vertical cavity surface emitting lasers (VCSELs) to provide a seamless grid-of-beam (GoB) coverage with multi-Gb/s per beam. We present systematic design and thorough analytical modeling of the AP architecture, which are then applied to downlink system modeling using non-imaging angle diversity receivers (ADRs). We propose static beam clustering with coordinated multi-beam joint transmission (CoMB-JT) for network interference management and devise various clustering strategies to address inter-beam interference (IBI) and inter-cluster interference (ICI). Non-orthogonal multiple access (NOMA) and orthogonal frequency division multiple access (OFDMA) schemes are also adopted to handle intra-cluster interference, and the resulting signal-to-interference-plus-noise ratio (SINR) and achievable data rate are derived. The network performance is studied in terms of spatial distributions and statistics of the downlink SINR and data rate through extensive computer simulations. The results demonstrate that data rates up to $15$ Gb/s are achieved within the coverage area and a properly devised clustering strikes a balance between the sum rate and fairness depending on the number of users.
@@ I Introduction
Optical wireless communication (OWC) technology based on narrow-beam infrared (IR) lasers has demonstrated the potential to provide ultra-high-speed transmissions for indoor multi-user wireless networks [3, 4, 5, 6, 7]. Beam-steered IR light communication is tailored to provide multi-Gb/s data rates per user by directing narrow IR laser beams at mobile devices [3, 8, 9]. One of the main challenges for the realization of indoor narrow-beam IR systems is beam steering as addressed in published research [10, 5, 6]. In [10], Koonen et al. have proposed and investigated a technique for two-dimensional (2D) steering of directional IR pencil beams (i.e., beams with very low divergence) by means of a passive diffractive optical module composed of crossed grating elements whose output beam angle is controlled by the wavelength of the laser beam. The authors have demonstrated downstream delivery with data rates of $32$ Gb/s and $42.8$ Gb/s per beam using $4$- pulse amplitude modulation (PAM) and adaptive discrete multitone (DMT) modulation, respectively, for a wavelength of $1550$ nm, with an angular coverage of $6^{\circ}\times 12^{\circ}$ at a $3$ m distance, equivalent to an area of $32\times 64$ cm2. In [5], Koonen et al. have proposed an alternative approach to implement wavelength-controlled 2D steering of IR pencil beams by using a high port-count arrayed waveguide grating router (AWGR) encompassing a bundle of single mode fibers with distinct wavelengths. To this end, they have rearranged the output fibers of the AWGR as a 2D square array and placed a lens next to the array, and the beam direction for each fiber is thereby determined by its position with respect to the optical axis of the lens. The authors have shown $112$ Gb/s $4$-PAM transmission based on an $80$-port C-band AWGR and a $9\times 9$ fiber array, aiming to cover an area of $0.75\times 0.75$ m2 with beam spots of diameter $8.5$ cm at a distance of $3.4$ m. They have then anticipated that this system can potentially achieve a total throughput of $8.9$ Tb/s (i.e., $80\times 112$ Gb/s). In [6], Wang et al. have designed a silicon photonic integrated phased array as an optical beam steering device, whereby they have achieved $12.5$ Gb/s error-free transmission over a distance of $1.4$ m using on-off keying (OOK) modulation.
@@
The aforementioned works apply advanced optical designs for beam steering to provide IR laser-class indoor coverage. However, the high complexity and expensive components associated with beam steering systems do not suit large-scale indoor network deployments. Taking a different approach, in [11, 12], Sun et al. have proposed a beam-domain massive multiple input multiple output (MIMO) optical wireless communication (OWC) system in which a large transmit lens is employed in front of a light-emitting diode (LED) array. The lens refracts lights emitted from the LED array elements toward different directions, enabling the optical access point (AP) to simultaneously communicate with a large number of user terminals in a wide coverage area. The authors have presented performance analysis for the optical massive multi-user MIMO transmission system with linear precoding and derived the optimal transmit covariance matrix under the total and per LED power constraints. Based on their findings, as the number of LEDs approaches infinity, the asymptotically optimal multi-user transmission policy under both power constraints is to use non-overlapping light beams, which is referred to as beam division multiple access (BDMA). According to [12], considering a scenario with four APs each made of a $12\times 12$ LED array and a transmit lens that cover $12$ users within a $5\times 5$ m2 cell in a room of height $3$ m, the data rate per user of $15$ bits per channel use is achieved based on BDMA for a $30$ dB transmitted signal-to-noise ratio (SNR).
Aiming for ultra-high transmission capacities and a full coverage for indoor environments, we have considered laser-based IR OWC systems by proposing an access network design in [13], and a backhaul system design in [14]. In both designs, we have used vertical cavity surface emitting laser (VCSEL) arrays to attain aggregate data rates beyond $1$ Tb/s. The choice of VCSELs over LEDs or other types of laser diodes is because of their prominent features including a high power efficiency, a high modulation bandwidth and well-controlled output beam properties. VCSELs are easier to fabricate and to be precisely arranged as 2D arrays in contrast to edge emitting lasers. They also have a low manufacturing cost, enhanced reliability, and circularly symmetric output beam profile [15]. In [16], Liu et al. have proposed an optical wireless transmitter system design based on a VCSEL array to provide a multi-beam coverage with uniform beam spots for indoor applications. The authors have adopted an optical design with multiple cascaded components for beam collimation, homogenization and expansion, by using micro-lens arrays, and plano-concave and plano-convex lenses. Based on a $5\times 5$ VCSEL array, they have demonstrated a total coverage of $1\times 1$ m2 with each beam spot of size $20\times 20$ cm2 in a $3$ m distance using Zemax OpticStudio simulations. As a proof of concept, the authors have also presented experimental results for a linear $1\times 4$ VCSEL array, confirming a data rate of $8$ Gb/s within each beam spot at a pre- forward error correction (FEC) bit error ratio (BER) of $3\times 10^{-3}$. They have anticipated that this transmitter design could achieve an aggregate data rate of $200$ Gb/s. However, the proposed structure is complex and requires meticulous design procedures to ensure a precise alignment between the optical components.
@@
In this paper, we propose a novel double-tier AP architecture incorporating an array of VCSEL arrays, which we also refer to as array of arrays for brevity. The AP architecture comprises multiple transmitter elements with each one having a predefined orientation to cover a specific area of the network. Each transmitter element is composed of a VCSEL array and a plano-convex lens. This design offers a rather simple and scalable solution suitable for establishing an indoor network of densely deployed laser beams to provide a seamless grid-of-beam (GoB) coverage through a single AP without the need for beam steering. To unlock the ultra-high capacity of the proposed AP and to eliminate strict alignment requirements, we employ a non-imaging angle diversity receiver (ADR) design based on compound parabolic concentrators coupled with high-speed photodiode (PD) arrays [17]. Besides, the contributions of this paper are summarized as follows:
@@
  • The multi-user performance of the GoB optical wireless network is thoroughly analyzed. To this end, the downlink system is elaborated in terms of the received signal-to-interference-plus-noise ratio (SINR) and achievable rate for both non-orthogonal multiple access (NOMA) and orthogonal frequency division multiple access (OFDMA) schemes, involving beam clustering and CoMB-JT in conjunction with non-imaging ADR and maximum ratio combining (MRC).
  • The system performance is evaluated through extensive computer simulations, providing key insights into the spatial GoB coverage and the impact of the distribution of user equipment (UE) devices in various beam clustering scenarios. The results evince how a judicious choice of the clustering layout besides the number of clusters can realize the best performance by maintaining the balance between the sum rate and fairness.
@@ II Gaussian Beam Propagation
A single-mode VCSEL generates the fundamental transverse electromagnetic (TEM) mode, designated as TEM00 mode, which is a spherical Gaussian beam with a circularly symmetric intensity profile. A Gaussian beam is characterized by the radius of the beam waist $w_{0}$ and the wavelength $\lambda$. The beam waist is where the wavefront is planar and the beam diameter is minimum. Assuming that the laser beam is propagating along the $z$ axis and the beam waist is located at the origin of cylindrical coordinates, the intensity distribution is expressed as [18, p. 84]:
@@
where $P_{\mathrm{t}}$ is the transmit optical power, and $r$ and $z$ are the radial and axial positions, respectively. The radius of the beam spot at distance $z$ from the beam waist is given by [18, p. 82]:
@@
By using (2), the divergence angle of the beam in the far field is obtained as [18, p. 85]:
@@
A Gaussian beam can also be fully characterized based on its complex $q$-parameter, which is defined as [18, p. 81]:
@@
In fact, the real and imaginary parts of the reciprocal of $q(z)$ encompass $R(z)$ and $w(z)$ as follows [18, p. 81]:
@@ III-A Array of Arrays Architecture
Fig. 1 depicts an indoor optical wireless network with full beam coverage by deploying a massive number of laser beams based on an array of arrays of VCSELs. The proposed double-tier AP design is composed of $9$ identical transmitter elements arranged as a $3\times 3$ array with a $5\times 5$ VCSEL array incorporated in each element [^1] The term ‘double-tier’ refers to the two-tier addressing process required to reach each VCSEL in the array of arrays architecture.. The total number of VCSELs used in this array of arrays architecture is $N_{\mathrm{VCSEL}}=225$. By choosing appropriate orientation angles for transmitter elements, each VCSEL array covers almost a ninth of the network area, as shown in Fig. 1. Note that the middle element of the AP that covers the central region of the network is not titled.
@@
Figs. LABEL:sub@Fig:AP_a and LABEL:sub@Fig:AP_b show the optical design of the double-tier AP. The adjacent transmitter elements are placed at a center-to-center distance of $d_{\mathrm{lens}}$. In each array, the VCSELs are placed close to each other with a pitch distance of $\delta$. As a result, the beam spots may significantly overlap on the receiver plane. The beams can be separated from one another by using a plano-convex lens in front of each VCSEL array. This allows the light beams to be to refracted towards different directions so as to provide a high spatial resolution on the receiver plane. The refraction angle experienced by each beam depends on the lens thickness and the path it travels through within the lens. In order to achieve the desired spatial resolution, the lens parameters need to be properly adjusted based on the design requirements. The adjustable parameters of the plano-convex lens include the diameter $L$, the radius of curvature $R_{\mathrm{lens}}$, and the center thickness $\tau_{\mathrm{c}}$ [19].
@@ III-B1 Light Refraction
When light traverses the boundary of two homogeneous media with refractive indices of $n_{1}$ and $n_{2}$, it is refracted according to Snell’s law so that [18, p. 7]:
@@
where $\theta_{1}$ is the angle of incidence, and $\theta_{2}$ is the angle of refraction, with respect to the normal line to the boundary. Snell’s law can be extended into a vector formulation to model the behavior of light rays as they traverse a curved surface separating two homogeneous media in 3D space. Sommerfeld and Runge demonstrated that this behavior can be analyzed using principles similar to those employed in electrodynamics for deriving the conditions that govern the changes of vector fields across a surface discontinuity [20]. Consequently, the law of refraction is derived analogously to the derivation of the continuity conditions for the tangential components of vector fields, leading to the vector form of Snell’s law [21, p. 133]:
$n_{1}\left(\mathbf{n}\times\mathbf{v}_{1}\right)=n_{2}\left(\mathbf{n}\times\mathbf{v}_{2}\right),$ (8)
or equivalently:
$\mathbf{n}\times\mathbf{v}_{2}=\mu\left(\mathbf{n}\times\mathbf{v}_{1}\right),\quad\mu=\frac{n_{1}}{n_{2}},$ (9)
where $\mathbf{n}$ is the unit vector normal to the boundary surface, $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ are the normalized incidence and refraction vectors, respectively, and $\times$ denotes the cross product. Note that the scaler form of Snell’s law is obtained by taking the Euclidean norm of both sides of (8). A key result is deduced from (8): the refracted ray lies in the same plane as the incident ray and the normal to the surface, which is referred to as the plane of incidence. The refraction vector $\mathbf{v}_{2}$ can be derived in terms of the incidence vector $\mathbf{v}_{1}$ and the normal vector $\mathbf{n}$ as follows:
@@ III-B2 Beam Transformation
$\mathbf{q}_{v}=\left[\begin{matrix}x_{v}\\ y_{v}\\ z_{v}\end{matrix}\right]=\mathbf{R}_{y}(\pi+\beta_{v})\mathbf{R}_{x}(\alpha_{v})\left[\begin{matrix}0\\ 0\\ d_{\mathrm{c}}\end{matrix}\right]+\mathbf{R}_{y}(\pi)\left[\begin{matrix}x^{\prime}_{v}\\ y^{\prime}_{v}\\ 0\end{matrix}\right]+\left[\begin{matrix}0\\ 0\\ h_{\mathrm{DL}}+d_{\mathrm{c}}\end{matrix}\right]=\left[\begin{matrix}-x^{\prime}_{v}-d_{\mathrm{c}}\sin\alpha_{v}\cos\beta_{v}\\ y^{\prime}_{v}-d_{\mathrm{c}}\sin\alpha_{v}\\ h_{\mathrm{DL}}+d_{\mathrm{c}}\left(1-\cos\alpha_{v}\cos\beta_{v}\right)\end{matrix}\right],$ (30)
In order to derive the parameters of the transformed beam, we use the $q$-parameter transformation [18, p. 97], [23]:
@@
By substituting the ABCD parameters from (20) into (22), the distance of the transformed beam waist from the tangent plane on the lens surface and the Rayleigh range of the transformed beam are derived as:
@@
where the operators $\mathrm{Re}$ and $\mathrm{Im}$ take the real and imaginary parts of a complex variable, respectively, and $f$ is the effective focal distance of the lens which, according to the lensmaker’s equation [24, p. 249], is defined as:
@@ III-C Total Spatial Intensity Distribution
In order to derive the total spatial intensity of the AP, the transmitter elements are indexed according to Fig. LABEL:sub@Fig:AP_b. Based on Euler angles with clockwise rotations, the tilting of the $v$th transmitter element about the $x$ and $y$ axes can be modeled using the following rotation matrices:
$\mathbf{R}_{x}(\alpha_{v})$ $=\left[\begin{matrix}1&0&0\\ 0&\cos\alpha_{v}&-\sin\alpha_{v}\\ 0&\sin\alpha_{v}&\cos\alpha_{v}\end{matrix}\right],$ (27a)
$\mathbf{R}_{y}(\beta_{v})$ $=\left[\begin{matrix}\cos\beta_{v}&0&\sin\beta_{v}\\ 0&1&0\\ -\sin\beta_{v}&0&\cos\beta_{v}\end{matrix}\right],$ (27b)
@@
where $\mathbf{v}^{\prime}_{i}$ is evaluated for $i=j-25(v-1)$ based on (17). The coordinates of the $v$th transmitter element are obtained as (30), given at the top of the page, where $h_{\mathrm{DL}}$ represents the vertical separation between the AP and the receiver plane, as shown in Fig. 1, $d_{\mathrm{c}}=d_{\mathrm{VL}}+\tau_{\mathrm{c}}$, and $\left[x^{\prime}_{v}~{}y^{\prime}_{v}\right]^{T}$ is equal to the $v$th column of $\mathbf{Q}_{\mathrm{t}}\in\mathbb{R}^{2\times 9}$ given by:
@@
(a) $R_{\mathrm{lens}}=25$ mm
(b) $R_{\mathrm{lens}}=17$ mm
(c) $R_{\mathrm{lens}}=15$ mm
@@
(a) $R_{\mathrm{lens}}=15$ mm
(b) $R_{\mathrm{lens}}=13$ mm
(c) $R_{\mathrm{lens}}=10$ mm